OpenStudy (anonymous):

x^3+27+0 please factor with steps

OpenStudy (smurfy14):

27+0? or is there another x?

OpenStudy (anonymous):

no his question is right

OpenStudy (anonymous):

(x+3)(x^2 -3x +9)

OpenStudy (anonymous):

+0 is kinda weird

OpenStudy (anonymous):

math textbooks are kinda weird. lol

OpenStudy (anonymous):

i meant =0 my bad

OpenStudy (anonymous):

(x+3)(x^2 -3x +9) = 0 x + 3 = 0 x = -3 x^2 -3x +9 = 0

OpenStudy (anonymous):

\[x^3+27=0\]\[\mathsf { We \ Can \ write \ it \ as \ (x)^3 + (3)^3 =0 } \]\[\mathsf {Now \ use \ the \ formula \ (x)^3 + (y)^3= (x+y)(x^2-xy+y^2)} \] \[\mathsf {(x+3)(x^2 -3x +9) = 0}\]\[\mathsf {x + 3 = 0 \ and \ x^2-3x+9=0 }\]\[\mathsf {x = -3} \] \[ \mathsf { x^2-3x+9=0 }\]